OpenStudy (anonymous):

-11=18+s/5 need a solution

OpenStudy (anonymous):

-55 = 18*5+s -55-18*5=s s = -145

OpenStudy (anonymous):

s/5 = -11-18=-29 x=5*(-29) x=-145

OpenStudy (saifoo.khan):

crrct

jimthompson5910 (jim_thompson5910):

-11 = 18 + s/5 -11*5 = 18*5 + s ... Note: multiply EVERY term by the LCD 5 to clear out the fraction -55 = 90 + s -55-90 = s -145 = s s = -145

OpenStudy (anonymous):

Got there first =P xD

OpenStudy (anonymous):

thanks guys and haha yea integral u did